A chemist needs 140 milliliters of a 33% solution but has only 11% AND 395 solutions available. Find how many milliliters of each that should be mixed to get the desired solution.
Answer by Alan3354(69443)

A chemist needs 140 milliliters of a 33% solution but has only 11% AND 395 solutions available. Find how many milliliters of each that should be mixed to get the desired solution.
------------
I think you mean 39%
---
e = amount of 11%
t = amount of 39%
---
e + t = 140
11e + 39t = 100*140
etc